OpenStudy (anonymous):

PLEASE HELP I CANT FIND THE ANSWER! Solve for x: 3(x + 1) = -2(x - 1) - 4. 1 -1 -5 -25

OpenStudy (uri):

@jewelshipp Use the rule of PEMDAS!!

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

you have to distribute the whole number into the polynomials first

OpenStudy (uri):

What was your final answer? @jewelshipp Please show your work.

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

i got like -2 and thats not an answer

OpenStudy (uri):

3(x + 1) = -2(x - 1) - 4. 3x+3=-2x+2-4 3x+2x=-2-3 5x=5 x=5/5 x=1

OpenStudy (uri):

-1*

OpenStudy (anonymous):

its negative 1

OpenStudy (anonymous):

not 1

OpenStudy (anonymous):

ok thanks

OpenStudy (uri):

@ilymartinez123 just corrected it.

OpenStudy (anonymous):

I know what i did wrong now thank you @uri

OpenStudy (uri):

:)